Let S be a finite nonempty set. Show that the number of subsets of S of even cardinality equals the number of subsets of S of odd cardinality.
1
Expert's answer
2020-11-17T16:51:18-0500

Let "|S|=n." In total there are "2^n" subsets of "S".


If "n" is odd then there is a one-to-one correspondence between sets with even cardinality and sets with odd cardinality. Subset "A" corresponds with its complement "S\setminus A". Consequently, the number of subsets with even cardinality equals the number of subsets with odd cardinality. So this number is "\frac{1}{2}2^n=2^{n-1}."


If "n" is even then put one element "a\in S" aside. With the trick described above we find that "S\setminus\{a\}"  has "2^{n-2}" subsets with even cardinality and also "2^{n-2}" subsets with odd cardinality. The subsets of "S\setminus\{a\}"  with odd cardinality become subsets "S"  with even cardinality if element "a" is added to each of them. This gives us "2^{n-2}+2^{n-2}=2^{n-1}" subsets of "S" with even cardinality. So in both cases the number of subsets of "S" of even cardinality equals the number of subsets of "S" of odd cardinality.